a) Calculate the increase in volume of 100cm^3 of murcury when its temperature changes from 10°C to 35°C. The volume coefficient of expansion of murcury, B is 0.00018°C^-1

b) Determine the change in volume of a block of cast iron 5.0cm × 10cm × 6.0cm, when the temperature changes from 15°C to 47°C. The coefficient of linear expansion, a for cast iron is 0.000010°C^-1.
1
Expert's answer
2018-08-28T09:44:08-0400
a)
The increase in volume of 100cm^3 of mercury is
ΔV=βΔTV
ΔV=(0.00018)(35-10)(100)=0.45 cm^3.

b)
The change in volume of a block of cast iron is
ΔV=3αΔTV
ΔV=3(0.000010)(47-15)(5)(6)(10)=0.29 cm^3.